OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 15 and the last term is -4,199,040? Answers to choose from. a)-3,599,175 b)-3,359,230 c)-3,119,285 d)-2,879,340

OpenStudy (campbell_st):

well you need to find the common ratio 1st the formula for a term in a geometric series is \[T_{n} = ar^{n -1}\] a = 1st term and r is the common ratio so you know \[-4199040 = 15\times r^{8 - 1}.... or -4199040 = 15 \times r^7\] you need to solve for r as the 1st step.

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

when you have found r the sum is \[S_{8} \frac{15(r^8 -1)}{r -1}\]

OpenStudy (anonymous):

i have that formula and for the answer i got -3,119,285 @campbell_st is this correct?

OpenStudy (campbell_st):

what did you get for r..?

OpenStudy (anonymous):

257

OpenStudy (campbell_st):

ummm big difference there this is how I found r \[\frac{-4199040}{15} = r^7\] so \[r = \sqrt[7]{\frac{-4199040}{15}}\] try that and see what you get

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

what did you get...?

OpenStudy (anonymous):

257 idk know whats wrong

OpenStudy (campbell_st):

ok... what is -4199040/15 =

OpenStudy (anonymous):

-279936

OpenStudy (campbell_st):

thats correct now on your calculator enter (-279936)^(1/7) fractional powers are the same as taking the 7th root.

OpenStudy (anonymous):

6

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

umm it should be r - 6

OpenStudy (campbell_st):

opps r = -6

OpenStudy (anonymous):

ohh ok my bad lol

OpenStudy (campbell_st):

ok... now you need to evaluate \[S_{8} = \frac{15((-6)^8 -1)}{(-6 -1)}\] you just need to evaluate the above... and the answer should be negative...and please input it as its written... otherwise you won't get the right answer.

OpenStudy (anonymous):

ok the answer i got this time is -2,879,340 is this correct?? @campbell_st

OpenStudy (campbell_st):

well i got -3599175 so you can choose.

OpenStudy (anonymous):

okay thank you for the guidance

OpenStudy (campbell_st):

glad to help